For a population with allele frequencies of f(M) = .2 and f(N) = .8, what are the EXPECTED genotypes under Hardy-Weinberg Equilibrium for the MN blood group? (hint: use the Hardy-Weinberg formula: p2 + 2pq + q2 = 1)
Answer #1

According to  Hardy-Weinberg Equilibrium genotypic frequency of the heterozygote is the 2*allelic frequency of M*allelic frequency of N

so frequency of MN blood group = 2*allelic frequency of M*allelic frequency of N

= 2*0.2*0.8

=0.32

so expected genotypic frequency of MN= 0.32

Expected genotypic frequency of MM= (allelic frequency of M)^2

= (0.2)^2

= 0.04

Expected genotypic frequency of NN= ( alleleic frequency of N)^2

= (0.8)^2

= 0.64